A machine cuts pieces of electromagnetic interference gaskets. The length of each gasket is measured and classified as being too short, OK, or too long. Likewise, the width of each gasket is measured and classified as being too narrow, OK, or too wide. The results of a sample of 1029 gaskets are shown in the table below. Test the claim that the width classification is independent of the length classification at the 90% confidence level.

Length

Width

Too Short

OK

Too Long

Totals

Too Narrow

10

98

7

115

OK

55

712

70

837

Too Wide

8

57

12

77

Totals

73

867

89

1029
Answer #1
Chi-Square Independence test

(1) Null and Alternative Hypotheses
The following null and alternative hypotheses need to be tested:
H0​: The two variables - Width classification and Length classification are independent
Ha​: The two variables - Width classification and Length classification are dependent

This corresponds to a Chi-Square test of independence.

(2) Degrees of Freedom
The number of degrees of freedom is df = (3 - 1) * (3 - 1) = 4

(3) Critical value and Rejection Region
Based on the information provided, the significance level is α=0.1, the number of degrees of freedom is df = (3 - 1) * (3 - 1) = 4, so the critical value is 7.7794.
Then the rejection region for this test becomes R={χ2:χ2>7.7794}.
1593201376410_image.png
(4)Test Statistics
The Chi-Squared statistic is computed as follows:
\chi^2 = \sum_{i=1}^n \frac{(O_{ij}-E_{ij})^2}{E_{ij}}=8.1854


(5)P-value
The corresponding p-value for the test is p=Pr(χ2​>8.1854)=0.085

(6)The decision about the null hypothesis
Since it is observed that χ2=8.1854>χ2_c​rit=7.7794, it is then concluded that the null hypothesis is rejected.

(7)Conclusion
It is concluded that the null hypothesis Ho is rejected. Therefore, there is enough evidence to claim that the two variables - Width classification and Length classification are dependent, at the 0.1 significance level.

Conditions:
a. The sampling method is simple random sampling.
b. The data in the cells should be counts/frequencies
c. The levels (or categories) of the variables are mutually exclusive.